# Comparing observed properties of winds in low-luminosity active galactic nuclei with theoretical predictions

### v1

@@ 1 Introduction
Both theory study and MHD numerical simulations of black hole accretion in the past decade have predicted the existence of strong wind launched from hot accretion flows (Blandford & Begelman, 1999; Yuan et al., 2012b, a; Narayan et al., 2012; Yuan et al., 2015; Bu et al., 2016). The detailed properties of wind, such as the velocity, mass flux, and special distribution have been investigated in detail based on three-dimensional GRMHD simulations (Yuan et al., 2015; Yang et al., 2021). Hot wind not only affects the dynamics of accretion inflow around black hole, but also serves as an important medium in AGN feedback (Fabian, 2012; Kormendy & Ho, 2013; Weinberger et al., 2017; Yuan et al., 2018; Yoon et al., 2019). By including the hot wind feedback in the numerical simulation study of the evolution of an individual elliptical galaxy, it has been shown that wind plays an important role in controlling the luminosity of black hole and its growth (Yoon et al., 2019).
Observational detection of wind driven by hot accretion flow, however, remains scarce and indirect for a long time. Wang et al. (2013) reported a flat rather than increasing density profile of hot accretion flow around Sgr A* from high-resolution X-ray spectroscopy, implying indirectly the existence of potential outflow suppress the inflow. Cheung et al. (2016) inferred the presence of centrally-driven, galactic-scale wind in a sample of typical quiescent galaxies hosting low-luminosity AGN around $z\sim 0.02$. The first direct observational evidence of hot wind driven by hot accretion flow in LLAGN has been detected through blue-shifted highly ionized iron emission lines in M81* and NGC 7213. The kinetic energy carried by these LLAGN winds can account for 10% – 15% of total AGN bolometric luminosity (Shi et al., 2021, 2022). The observational evidence for the interaction between wind and the cicumnuclear gas within pc-scale has been found and is likely the physical mechanism of keeping the central AGN dim (Shi et al., 2024).
@@ 2.2 Sample Selection
The bolometric luminosity of AGN implies the energy released by the accretion material at the event horizon of the central black hole. The mass inflow rate at the event horizon would be estimated by $\dot{M}_{\rm in}=\frac{L_{\rm bol}}{\epsilon c^{2}}$, where $\epsilon$ refers to the radiation efficiency. For radiatively inefficient hot accretion flows, $\epsilon$ decreases with the decreasing of the Eddington ratio $R_{\rm Edd}\equiv\frac{L_{\rm bol}}{L_{\rm Edd}}$ and falls below that of the standard thin disk ($\epsilon_{\rm SSD}\simeq 0.1$) in the low-luminosity regime. In this paper, we adopt the radiation efficiency $\epsilon$ from the theoretical $\epsilon$$\frac{\dot{M}_{\rm in}}{\dot{M}_{\rm Edd}}$ correlation provided in Xie & Yuan (2012), assuming the fraction from viscous heating is $\delta\sim 0.5$. The Eddington accretion rate is defined as $\dot{M}_{\rm Edd}\equiv 10L_{\rm Edd}/c^{2}$. So the ratio between the net inflow rate at event horizon and the Eddington accretion rate are linked to the Eddington ratio $R_{\rm Edd}\equiv\frac{\epsilon\dot{M}_{\rm in}}{0.1\dot{M}_{\rm Edd}}$. Thus given any Eddington ratio $R_{\rm edd}$ we would be able to precisely determine the net inflow rate at event horizon.
@@ 3 Results
The right panel of Fig. 1 shows the comparison between $r_{\rm tr}$ and wind location $R_{\rm max}$ deduced from photoionization modeling. We can see from the figure that, the estimated $r_{\rm tr}$ for both UFO and hot wind are in line with $R_{\rm max}$, especially in the case of MAD with high BH spin. In contrast, WAs lie above the line where $R_{\rm max}=r_{\rm tr}$. For UFO and hot wind, the detection location is close to the black hole and more specifically, close to the truncation radius. So the detected wind velocity would be roughly consistent with the initial velocity of wind at its launching point. Our assumption is valid that the wind velocity originated from truncation radius remains almost constant before encountering ISM (Cui & Yuan, 2020). But for WA, the detection location seems far away from the wind launching region. One possible scenario is that although in LLAGNs outflows are driven mainly from $r_{\rm tr}$, the detected WAs trace the wind components located at a larger radius ($\gtrsim 10^{5}\rm~{}r_{g}$) away from their origin points. Only at such large distances, the ionization state of WA is lower and “suitable” to be detected in the soft X-rays. As the distance is further away, WAs may be opt to experience deceleration through interacting with interstellar medium (ISM), resulting in a larger putative truncation radius according to Eq. (2) - (5) with smaller wind velocity than expected. Since WAs may not trace the launching site of winds in the LLAGN, the intrinsic truncation radius of these AGN would be much smaller than deduced. And the mass outflow rate estimation of these WAs would not reflect the intrinsic black hole wind.
Wind driven by hot accretion flow is expected to carry a substantial amount of material away. The radial profile of its mass outflow rate can be described by $\dot{M}_{\rm out,max}=\dot{M}_{\rm in}(\frac{r}{40\rm~{}r_{g}})$ according to numerical simulations (Yuan et al., 2015). $\dot{M}_{\rm in}$ refers to the mass accretion rate at event horizon (See in Sec. 2.2). We investigate the ratio between the wind mass flux and black hole accretion rate for our sample. The results are shown in Figure 2. For UFO together with hot wind, the ratio is consistent with the theoretical estimates, while all the WAs are systematically off-set from the expected correlation.

### v2

@@ 1 Introduction
Both theory study and MHD numerical simulations of black hole accretion in the past decade have predicted the existence of strong wind launched from hot accretion flows (Blandford & Begelman, 1999; Yuan et al., 2012b, a; Narayan et al., 2012; Yuan et al., 2015; Bu et al., 2016). The detailed properties of the hot wind, such as the velocity, mass flux, and special distribution have been investigated based on three-dimensional GRMHD simulations(Yuan et al., 2015; Yang et al., 2021). Hot wind not only affects the dynamics of accretion inflow around black hole, but also serves as an important medium in AGN feedback (Fabian, 2012; Kormendy & Ho, 2013; Weinberger et al., 2017; Yuan et al., 2018; Yoon et al., 2019). By including the hot wind feedback in the numerical simulation study of the evolution of an individual elliptical galaxy, it has been shown that wind plays an important role in controlling the luminosity of black hole and its growth (Yoon et al., 2019).
Observational detection of wind driven by hot accretion flow, however, remains scarce and indirect for a long time. Wang et al. (2013) reported a flat rather than increasing density profile of hot accretion flow around Sgr A* from high-resolution X-ray spectroscopy, implying indirectly the existence of potential outflow suppress the inflow. Cheung et al. (2016) inferred the presence of centrally-driven, galactic-scale wind in a sample of typical quiescent galaxies hosting low-luminosity AGN around $z\sim 0.02$. The first direct observational evidence of hot wind driven by hot accretion flow in LLAGN has been detected through blue-shifted highly ionized iron emission lines in M81* and NGC 7213. The kinetic energy carried by these LLAGN winds can account for 10% – 15% of total AGN bolometric luminosity (Shi et al., 2021, 2022). This is larger than the theoretical prediction that the mechanical efficiency generated by the hot accretion flows can hardly exceed 3% of the accretion mass energy flux claimed by Sądowski & Gaspari (2017). The observational evidence for the interaction between wind and the cicumnuclear gas within pc-scale has been found and is likely the physical mechanism of keeping the central AGN dim (Shi et al., 2024).
@@ 2.2 Sample Selection
The bolometric luminosity of AGN implies the energy released by the accretion material at the event horizon of the central black hole. The mass inflow rate at the event horizon would be estimated by $\dot{M}_{\rm in}=\frac{L_{\rm bol}}{\epsilon c^{2}}$, where $\epsilon$ refers to the radiation efficiency. For radiatively inefficient hot accretion flows, $\epsilon$ decreases with the decreasing of the Eddington ratio $R_{\rm Edd}\equiv\frac{L_{\rm bol}}{L_{\rm Edd}}$ and falls below that of the standard thin disk ($\epsilon_{\rm SSD}\simeq 0.1$) in the low-luminosity regime. In this paper, we adopt the radiation efficiency $\epsilon$ from the theoretical $\epsilon$$\frac{\dot{M}_{\rm in}}{\dot{M}_{\rm Edd}}$ correlation provided in Xie & Yuan (2012), assuming the fraction from viscous heating is $\delta\sim 0.5$. Their results are consistent with Sądowski & Gaspari (2017). The Eddington accretion rate is defined as $\dot{M}_{\rm Edd}\equiv 10L_{\rm Edd}/c^{2}$. So the ratio between the net inflow rate at event horizon and the Eddington accretion rate are linked to the Eddington ratio $R_{\rm Edd}\equiv\frac{\epsilon\dot{M}_{\rm in}}{0.1\dot{M}_{\rm Edd}}$. Thus given any Eddington ratio $R_{\rm edd}$ we would be able to precisely determine the net inflow rate at event horizon.
@@ 3 Results
The right panel of Fig. 1 shows the comparison between $r_{\rm tr}$ and wind location $R_{\rm max}$ deduced from photoionization modeling. We can see from the figure that, the estimated $r_{\rm tr}$ for both UFO and hot wind are in line with $R_{\rm max}$, especially in the case of MAD with high BH spin. In contrast, WAs lie above the line where $R_{\rm max}=r_{\rm tr}$. For UFO and hot wind, the detection location is close to the black hole and more specifically, close to the truncation radius. So the detected wind velocity would be roughly consistent with the initial velocity of wind at its launching point. Our assumption is valid that the wind velocity originated from truncation radius remains almost constant before encountering ISM (Cui & Yuan, 2020). But for WA, the detection location seems far away from the wind launching region. One possible scenario is that although in LLAGNs outflows are driven mainly from $r_{\rm tr}$, the detected WAs trace the wind components located at a larger radius ($\gtrsim 10^{5}\rm~{}r_{g}$) away from their origin points. Only at such large distances, the ionization state of WA is lower and “suitable” to be detected in the soft X-rays. As the distance is further away, WAs may be opt to experience deceleration through interacting with interstellar medium (ISM), resulting in a larger putative truncation radius according to Eq. (2) - (5) with smaller wind velocity than expected. Since WAs may not trace the launching site of winds in the LLAGN, the intrinsic truncation radius of these AGN would be much smaller than deduced. And as the distance increases, the initially driven wind is expected to entrain more and more mass(Gaspari & Sądowski, 2017). So the mass outflow rate estimation of these WAs would not reflect the intrinsic black hole wind.
Wind driven by hot accretion flow is expected to carry a substantial amount of material away. The radial profile of its mass outflow rate can be described by $\dot{M}_{\rm out,max}=\dot{M}_{\rm in}(\frac{r}{40\rm~{}r_{g}})$ according to numerical simulations (Yuan et al., 2015). $\dot{M}_{\rm in}$ refers to the mass accretion rate at event horizon (See in Sec. 2.2). We investigate the ratio between the wind mass flux and black hole accretion rate (also known as the mass loading factor) for our sample. The results are shown in Figure 2. For UFO together with hot wind, the ratio is consistent with the theoretical estimation, while all the WAs are systematically offset from the expected correlation. Fiore et al. (2024) collects a sample of wind components mostly detected in AGN with intermediate and high accretion rate and suggests a negative power-law correlation between the mass loading factor and the Eddington ratio. However in our LLAGN wind sample, we do not find such a strong correlation and the mass loading factor are on average larger than their high accretion rate counterparts. This adds further support to a different wind launching mechanism between hot mode and cold mode accretion.